COMPLEX CROWNS OF RIEMANNIAN SYMMETRIC SPACES AND NON-COMPACTLY CAUSAL SYMMETRIC SPACES

SIMON GINDIKIN BERNHARD KRÖTZ

Abstract. In this paper we define a distinguished boundary for the complex crowns Ξ ⊆ G_{ℂ}/K_{ℂ} of non-compact Riemannian symmetric spaces G/K. The basic result is that affine symmetric spaces of G can appear as a component of this boundary if and only if they are non-compactly causal symmetric spaces.